OpenStudy (anonymous):

How to find the roots to a equation with complex coefficients? I tried one, but i'm stuck

OpenStudy (anonymous):

\[2x^2-ix-2+i\]

OpenStudy (phi):

You can use the quadratic formula with a=2, b= -i and c= (-2+i) But you will have to take the square root of a complex number. See https://en.wikipedia.org/wiki/Square_root#Algebraic_formula for one way to do this.

OpenStudy (anonymous):

I did it..but I got \[x^4+17x^2-16\]

OpenStudy (anonymous):

and idk how to solve that

OpenStudy (anonymous):

my prof does it differently

OpenStudy (anonymous):

and i dont get his way

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\[x=\frac{ \iota \pm \sqrt{\left( -\iota \right)^{2}-4*2\left( -2+\iota \right)} }{ 2*2 }\] \[=\frac{ \iota \pm \sqrt{-1+16-8\iota } }{ 4 }\] \[=\frac{ \iota \pm \sqrt{15-8\iota} }{4 }=\frac{ \iota \pm \sqrt{16-1-8\iota} }{ 4 }\] \[=\frac{ \iota \pm \sqrt{\left( 4 \right)^{2}+\iota ^{2}-2*4*\iota} }{4 }\] \[=\frac{ \iota \pm \left( 4-\iota \right) }{4 }\] now you can solve to find the two values.

OpenStudy (anonymous):

omg..nvm..i switched the -ive in the -4 part of the quadratic formula to a +

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

can u help me with something

OpenStudy (anonymous):

I'm getting the wrong roots

OpenStudy (anonymous):

can someone check what I'm doing wrong

OpenStudy (amoodarya):

hint : use a=x^2

OpenStudy (amoodarya):

then you will have a^2+17a-16=0 (a+1)(a-16)=0 a=-1 ,a=16 x^2=-1 ---> x=+-i x^2=16 --->x=+-4

OpenStudy (amoodarya):

or from the first it seems x^2=-1 is a factor check it or see below x^2=-1 x^4+17x^2-16 =(-1)^2+17(-1)-16=0 so you can divide x^4+17x^2-16 by x^2+1

OpenStudy (amoodarya):

2x^2−ix−2+i =2x^2-2−ix+i =2(x^2-1) -i(x-1) now you can factor (x-1)

OpenStudy (anonymous):

ok. when i put it into the quadratic form and try to find the roots. Do i use all 4 of the roots that i get

OpenStudy (anonymous):

could i post my work and you check it over?

OpenStudy (amoodarya):

\[2x^2-ix-2+i =0 \rightarrow \Delta=\sqrt{i^2-4(2)(-2+i)}\] \[\Delta=\sqrt{i^2-4(2)(-2+i)} =\sqrt{-1-8(-2+i)}=\sqrt{-1+16-8i}\] \[\sqrt{-1+16-8i}=\sqrt{(4-i)^2}=4-i\]

OpenStudy (anonymous):

i posted my work

OpenStudy (amoodarya):

\[x =\frac{ +i \pm \Delta }{ 4 }=\frac{ +i \pm(4-i) }{ 4 }\]

OpenStudy (anonymous):

ok..i looked up a video and it said that i only use 1 of the sets of roots

OpenStudy (anonymous):

could u look at my work and correct me where im going wrong

OpenStudy (amoodarya):

it was very unclear but i find your mistake (-1+16-8i)=(4-i)^2 not (-1+4i)^2

OpenStudy (amoodarya):

(4-i)^2=16+i^2-8i but (-1+4i)^2=1+16i^2-8i=1-16=8i

OpenStudy (anonymous):

how did u get (4-i)

OpenStudy (amoodarya):

-1+16 -8i =i^2 +4 ^2 -2 *4 * i

OpenStudy (anonymous):

???

OpenStudy (anonymous):

could u write it using the equation function on here

OpenStudy (amoodarya):

function ! it was (a-b)^2 =a^2 +b^2 -2 a b just!

OpenStudy (anonymous):

I get 15+8i...and then i evaluate it on the far right side of my work....

OpenStudy (anonymous):

can u look there and see what im doing wrong

OpenStudy (amoodarya):

ok begin with 15+8i \[15+8i =\sqrt{15^2+8^2} e^{i \theta}\] \[\sqrt{225+64} e^{i \theta}=17 e^{i \theta} \] \[\theta = \tan ^{-1} \frac {y}{x} =\tan ^{-1} \frac {8}{15} \] 

OpenStudy (anonymous):

can u not use e..i didnt learn it

OpenStudy (amoodarya):